# Remote Sensing Reveals Adoption of Sustainable Rice Farming Practices Across Punjab, India

### v1

@@ Appendix C Modeling Details
We experiment with ensemble-based learning methods including random forest (RF) and gradient-boosted trees (GB) to classify rice plots into one of 3 possible classes: AWD, DSR or PTR. Previous works have shown the suitability of these methods to LULC and agricultural practice mappings [32, 33, 40, 41]. Since we can reduce plot-level features to a point estimate, time-series modeling is possible with ensemble models like RF or GB. We use scikit-learn’s implementation for RF and LightGBM for GBs.
@@
Hand-crafted Features: We generate a set of hand-crafted features from each S-1 band ($VV,VH,VV/VH$) per orbit, as the summary statistics for a rice-growing plot. For each band, we locate the positions of the first $K$ troughs and crests as well as the first $K$ inflection points. We extract the timestamp and amplitude at these points, measuring timestamp relative to the first date in the timeseries, in number of days. Experimentally, it was found that $K=3$ would cover all timeseries in the training dataset, with larger $K$s leading to missing values. We also featurize the total number of trough and crests, along with the mean, minimum and maximum amplitude values of each band per orbit. Following previous work [28], we fit a Gaussian kernel to the VV/VH ratio and employ its parameters (amplitude, day of peak, standard deviation and $R^{2}$ of kernel fit) as additional features. Additionally, we calculate the Radar Vegetation Index (RVI) as modified for S-1[30] as it has been shown to be useful in detecting crop height. The timeseries generation process for S-1 bands after processing and additional hand-crafted features are visualized in Figure A2.
@@
Presto Features: In addition to hand-crafted features, we employ the Pretrained Remote Sensing Transformer (Presto) [31] as a temporal feature extractor. Presto is transformer-based deep learning model that has been pre-trained on remote sensing pixel-timeseries data, including Sentinel-1 SAR, Sentinel-2 multispectral, Dynamic Earth land use and land cover (LULC) labels, ERA5 temperature and precipitation datasets and, latitude & longitude. To extract embeddings, we freeze the Presto encoder, fix the Dynamic Earth class to ‘cropland’, and input both ERA5 variables and Sentinel-1 VV and VH bands from either ascending or descending orbits. Since we fit a spline to smoothed SAR timeseries, we can use arbitrary sampling frequencies to pseudo-assess the effect of sampling frequency. This is especially important as a limitation of the Presto encoder is that it cannot process more than 24 timesteps-in such cases we can increase the sampling duration to fit the entire timeseries. Therefore we choose $f$=4,7,10 days. Additionally, we extract embeddings from the encoder at each timestep  [^7] Presto takes a timestamp numbering for the month number, however we find that it gives better performance to use the week number instead of the month, since irrigation and sowing signal frequency is higher than a month, plus one feature for the location, generating a total of $1128\times(T\times 2)+$ 1888One set of embeddings for S1 and Dynamic World inputs embeddings. This process is encapsulated within the ‘Presto’ block of Fig. A6.
C.1 Training Pipeline
The full training pipeline is shown in Figure A6.
@@ C.1 Training Pipeline
Fig. A3: Impact of orbit direction and climatology data (temperature and precipitation) on overall accuracy of DSR and AWD classification performance. Left: Bar plot grouped by orbit direction (ascending, descending, both) with confidence intervals highlighting variability within each task. Right: Bar plot showing the impact of climatology inclusion (yes, no) on OA, with confidence intervals highlighting variability within each task.
C.2 Inference Pipeline
The full inference pipeline is shown in Figure A7.
Appendix D Ablations
D.1 Effect of Sentinel-1 Orbit
@@
For DSR classification, ascending orbits consistently outperformed descending orbits, as shown in Figure A3 (left). This suggests that radiometric properties of the descending orbit make it suboptimal for capturing DSR features in this region. The use of both orbits showed no significant advantage over ascending orbit alone, and the additional effort required to process both orbits may not be justified. For AWD classification, orbit direction had less impact, likely due to lag features driving performance regardless of orbit, while AWD-specific features were not captured effectively. This highlights that ascending orbits better capture DSR features compared to descending orbits in this study.
D.2 Effect of Climatology Data
@@
High overlap in CIs for models with and without climatology data indicates no significant benefit from including these variables (Figure A3 , right). As the average performance without ERA5-Land data was slightly higher, and considering the effort required to acquire and integrate this dataset, there appears to be no inherent advantage to its inclusion for this region and task.
@@
Appendix E Error Analysis
On the held out test set of 129 plots from our stratified split, we find that for DSR classification,
@@ Appendix E Error Analysis
Given the distributed but nearly separable cropping calendar (Figure A1), it is useful to analyze the sources of classification errors with respect to the original classes. For the best-performing DSR classification model, Figure A4 (top-left) highlights the proportion of test set errors originating from each original class, showing a linear decrease in errors as planting dates increment from left to right. Similarly, Figure A4 (top-right) illustrates the proportion of incorrect samples and their original classes for AWD classification, with errors decreasing towards the left, for dates further from the AWD class.
@@
For both tasks, errors predominantly stem from temporally similar samples (~33% of that class if incorrectly classified), suggesting that lag features may not provide sufficient discrimination in these cases and that practice features lack the necessary power. Confusion matrices (Fig. A4 bottom row) reveal that negative class scores dominate the results. For both classification tasks, the positive classes(DSR, AWD) achieves precision of 75% and recall of 67%.
Fig. A4: Error Analysis. Top-left: DSR classification error percentages on test set from original classes. Bottom-left: DSR classification confusion matrix from collapsed classes. Top-right: AWD classification error percentages on test set from original classes. Bottom-right: AWD classification confusion matrix from collapsed classes. Text on bars represents percentage of errors contributed from that class.
E.1 Feature Importance
@@
Fig. A5: Feature importance: Using the gain attributed to each feature, the feature importance for LightGBM models are estimated for sowing (DSR) classification, left and irrigation (AWD) classification on the right.
@@
For the best performing models for DSR and AWD classification, we estimate the feature importance as shown in Figure A5. For DSR classification, the best performing model was trained on the time period Jun 1- Sep 15 with a 4 day sampling frequency (Figure 4, row 3). Therefore the features that start with “time_*” refer to the time period in 4 day increments beginning on Jun 1999Since we pass in both S1 timesteps and a single fixed value for Dynamic World, the Presto Encoder generates: [0] lat-lon token, [1:T+1] temporal tokens, [T+1:2T+1] Dynamic World tokens. The top two features are therefore from the time period Jun 16-20, implying that the features in those regions were most useful in classification. From Figure A1, using the lens of sowing detection, we see that this refers to the time when the majority of PTR crops have been sowed (control/DSR has peaked and AWD has begun), which aligns with our conclusion that the sowing features are actually helpful.
@@
For AWD classification the model was trained for the period May 1 to Dec 15 with a ten day sampling frequency (Figure 4, row 12). Similarly, the top two features (Figure A5 left) correspond to the time periods of June 20-30, and Sep 18-28. The former is roughly the same timeperiod as the top-2 features for DSR, while the latter likely captures some important lag feature, since all plots are at some stage of growth in September, regardless of when they were planted.
Appendix F Large-scale Inference
@@
Thereafter, we follow the method laid out in subsection 4.5 with ascending orbit only, and use the best pretrained model for DSR classification, without any climatology inputs. The combination of pretrained models such as Presto and FTW, along with our dimensional training methodology that requires no cropping calendar information, allows us to apply this methodology to arbitrary spatial scales.
Fig. A6: Training and validation workflow used in the study. We utilize Copernicus Dataspace and GEE to acquire Sentinel-1 and ERA5 temporal data on a per plot basis, and to subsequently reduce them to plot level statistics. We then fit a spline, resample, generate hand-crafted features and embeddings from Presto over a given time period. The separability of water management classes is modeled using a Random Forest and Gradient Boosted Trees. The inset shows an example of the average S-1 VV and VH timeseries after smoothing and spline fitting, with confidence intervals at each point in time.
Fig. A7: Inference Workflow: Methodology followed during DSR classification predictions for 3 million polygons using S-1 timeseries. FTW[20] polygons are generated from bitemporal 4-bands S-2 imagery

### v2

@@ Appendix C Modeling Details
Sentinel-1 timeseries is generated by the process explained in section 3. Additionally, for all features, we experiment with resampling the smoothed spline at frequencies of 4, 7 and 10 days. This allows us to modulate the temporal resolution of the data where possible (when the total number of timesteps $\leq$ 24) or to accommodate longer temporal ranges (such as rows 11-12 in Table 2).
@@
We experiment with ensemble-based learning methods including random forest (RF) and gradient-boosted trees (GB) to classify rice plots into one of 3 possible classes: AWD, DSR or PTR. Previous works have shown the suitability of these methods to LULC and agricultural practice mappings  (liuUsingSentinel1Sentinel22022; fangComprehensiveReviewRice2024a; rudiyantoAutomatedNearRealTimeMapping2019; TATSUMI2015171). Since we can reduce plot-level features to a point estimate, time-series modeling is possible with ensemble models like RF or GB. We use scikit-learn’s implementation for RF and LightGBM for GBs. Additionally we employ Dynamic Time Warping along with kNN classification as baseline measure of performance, that directly compares each test timeseries to the training datasets to infer class.
Hand-crafted Features
@@
We generate a set of hand-crafted features from each S-1 band ($V$,VH,VV/VHVV,VH,VV/VH) per orbit, as the summary statistics for a rice-growing plot. For each band, we locate the positions of the first $K$ troughs and crests as well as the first $K$ inflection points. We extract the timestamp and amplitude at these points, measuring timestamp relative to the first date in the timeseries, in number of days. Experimentally, it was found that $K=3$ would cover all timeseries in the training dataset, with larger $K$s leading to missing values. We also featurize the total number of trough and crests, along with the mean, minimum and maximum amplitude values of each band per orbit. Following previous work (bazziMappingPaddyRice2019), we fit a Gaussian kernel to the VV/VH ratio and employ its parameters (amplitude, day of peak, standard deviation and $R^{2}$ of kernel fit) as additional features. Additionally, we calculate the Radar Vegetation Index (RVI) as modified for S-1 (nasirzadehdizaji2019sensitivity) as it has been shown to be useful in detecting crop height. The timeseries generation process for S-1 bands after processing and additional hand-crafted features are visualized in Fig.  A11.
Presto Features
@@
In addition to hand-crafted features, we employ the Pretrained Remote Sensing Transformer (Presto) (tseng2023lightweight) as a temporal feature extractor. Presto is transformer-based deep learning model that has been pre-trained on remote sensing pixel-timeseries data, including Sentinel-1 SAR, Sentinel-2 multispectral, Dynamic Earth land use and land cover (LULC) labels, ERA5 temperature and precipitation datasets and, latitude & longitude. To extract embeddings, we freeze the Presto encoder, fix the Dynamic Earth class to ‘cropland’, and input both ERA5 variables and Sentinel-1 VV and VH bands from either ascending or descending orbits. Since we fit a spline to smoothed SAR timeseries, we can use arbitrary sampling frequencies to pseudo-assess the effect of sampling frequency. This is especially important as a limitation of the Presto encoder is that it cannot process more than 24 timesteps-in such cases we can increase the sampling duration to fit the entire timeseries. Therefore we choose ff=4,7,10 days. Additionally, we extract embeddings from the encoder at each timestep  [^6] Presto takes a timestamp numbering for the month number, however we find that it gives better performance to use the week number instead of the month, since irrigation and sowing signal frequency is higher than a month, plus one feature for the location, generating a total of $128\times(T\times 2)+1$ [^7] One set of embeddings for S1 and Dynamic World inputs embeddings. This process is encapsulated within the ‘Presto’ block of Fig. A3.
AlphaEarth Embeddings
AlphaEarth Foundations (brown2025alphaearth) generates geospatial embeddings through a spatiotemporal encoder–decoder architecture that jointly models multi-source Earth observation inputs across continuous time. Each embedding represents a 64-byte vector summarizing environmental and surface dynamics over a defined “valid period”, learned from heterogeneous inputs (e.g., Sentinel-1/2, Landsat, GEDI, ERA5-Land, GRACE, NLCD, and Wikipedia text). The model’s Space-Time Precision (STP) encoder integrates spatial self-attention, temporal axial attention, and convolutional “precision” operators with Laplacian pyramid rescaling to preserve both global context and local spatial detail. An implicit decoder reconstructs sensor-specific variables from the embeddings using conditional time and sensor metadata, regularized by a batch-uniformity loss enforcing uniform distribution on the unit hypersphere $S^{63}$. Training employs contrastive learning between teacher–student video models and a text-alignment branch, producing dense, continuous 10 m embeddings that encode the joint spatiotemporal structure of the Earth system, which is what we use as the AE product, available from Google Earth Engine at this link: https://developers.google.com/earth-engine/datasets/catalog/GOOGLE˙SATELLITE˙EMBEDDING˙V1˙ANNUAL.
Modeling Details
For RF, we sweep over the following ranges for each hyperparameter:
  • n_estimators: [300, 1800]
  • max_depth: [2, 20]
  • min_samples_split: [2, 20]
  • min_samples_leaf: [1,20]
For GB, we sweep over the following ranges for each hyperparameter:
  • max_depth: [3,8]
  • learning_rate [0.001, 0.01]
  • n_estimators: [100, 500]
  • subsample: [0.6,1.]
  • colsample_bytree: [0.6, 1.]
  • min_child_samples: [5, 30]
  • min_split_gain: [1e-8, 0.1]
  • num_leaves: [20, 3000]
  • reg_alpha: [1e-8, 0.5]
  • reg_lambda: [1e-8, 0.5]
All model details can be found in the code file under scripts/train_model.py
For all models and feature combinations, except the random baseline, we run a single training run with a seed=42. The random baseline is run 1000 times and the mean±stdevs are reported in  Table 1.
All model training was done on a x86_64 Linux machine with 250Gb of RAM and 64 CPUs. A L40s GPU was used to generate the Presto embeddings, but this can be achieved on any CPU. All software versions are available in requirements.txt file.
Final model weights for the best DSR and AWD models are also provided in data/models.
@@
Figure A4: Impact of orbit direction and climatology data (temperature and precipitation) on overall accuracy of DSR and AWD classification performance. Left: Bar plot grouped by orbit direction (ascending, descending, both) with confidence intervals highlighting variability within each task. Right: Bar plot showing the impact of climatology inclusion (yes, no) on OA, with confidence intervals highlighting variability within each task.
Figure A5: Error Analysis. Top-left: DSR classification error percentages on test set from original classes. Bottom-left: DSR classification confusion matrix from collapsed classes. Top-right: AWD classification error percentages on test set from original classes. Bottom-right: AWD classification confusion matrix from collapsed classes. Text on bars represents percentage of errors contributed from that class.
Appendix D Foundation Model Performance
From table 1, it is evident that features derived from remote-sensing foundation models–Presto and AlphaEarth (AE)–consistently dominate performance, particularly when combined with complementary feature families such as Fourier (FT) and handcrafted (HC) descriptors. While HC features require domain expertise in SAR signal behavior and agronomic processes to construct, FT and DTW are standard numerical transformations applicable to most temporal datasets with minimal preprocessing.
The strong performance of FT–FM combinations underscores a key finding: foundation model embeddings capture rich temporal representations that align well with frequency-domain features, jointly enhancing separability of both sowing and irrigation practices. This synergy suggests a practical route for operational deployments, especially for practitioners lacking extensive SAR expertise.
Among the two models, AE offers the most accessible path to adoption, as its embeddings are freely available on Google Earth Engine for research and non-profit use. In contrast, Presto requires local preprocessing of Sentinel-1 data (noise removal, calibration, speckle filtering, terrain correction) before embeddings can be generated, though it remains valuable when high temporal fidelity is required. Together, these models highlight the growing utility of pretrained Earth-observation representations for agricultural monitoring: practitioners can directly extract embeddings, pair them with simple numerical transforms such as FT, and train lightweight classifiers for diverse crop-mapping tasks without retraining large models or acquiring new ground data.
Surprisingly, AE embeddings—which summarize temporal dynamics over a full calendar year—achieved the highest performance, despite rice cultivation in Punjab occurring only between April and December and varying by several months across plots. This suggests that AE captures persistent spectral–textural cues linked to sowing and irrigation transitions, even when embedded within multi-crop annual cycles. Further investigation is needed to understand why features derived from year-long multisensor inputs remain sensitive to rice-specific signals without explicit temporal masking, unlike Presto, for which we constrained inputs to the active growing season.
Appendix E Ablations
Effect of Sentinel-1 Orbit
@@ Appendix E Ablations
For DSR classification, ascending orbits consistently outperformed descending orbits, as shown in Fig. A5 (left). This suggests that radiometric properties of the descending orbit make it suboptimal for capturing DSR features in this region. The use of both orbits showed no significant advantage over ascending orbit alone, and the additional effort required to process both orbits may not be justified. For AWD classification, orbit direction had less impact, likely due to lag features driving performance regardless of orbit, while AWD-specific features were not captured effectively. This highlights that ascending orbits better capture DSR features compared to descending orbits in this study.
Effect of Climatology Data
@@
High overlap in CIs for models with and without climatology data indicates no significant benefit from including these variables (Fig. A5 , right). As the average performance without ERA5-Land data was slightly higher, and considering the effort required to acquire and integrate this dataset, there appears to be no inherent advantage to its inclusion for this region and task.
@@
Appendix F Error Analysis
@@
Given the distributed but nearly separable cropping calendar (Fig. A1), it is useful to analyze the sources of classification errors with respect to the original classes. For the best-performing DSR classification model, Fig. A5 (top-left) highlights the proportion of test set errors originating from each original class, showing a linear decrease in errors as planting dates increment from left to right. Similarly, Fig. A5 (top-right) illustrates the proportion of incorrect samples and their original classes for AWD classification, with errors decreasing towards the left, for dates further from the AWD class.
@@
For both tasks, errors predominantly stem from temporally similar samples ($\sim$33% of that class if incorrectly classified), suggesting that lag features may not provide sufficient discrimination in these cases and that practice features lack the necessary power. Confusion matrices (Fig. A5 bottom row) reveal that negative class scores dominate the results. For both classification tasks, the positive classes(DSR, AWD) achieves precision of 75% and recall of 67%.
Importance of orbit direction:
For operational monitoring at scale, we found ascending orbit data consistently outperforms descending orbit data for DSR classification, with no benefit from combining orbits or including climatology variables (Figure A5). This streamlined approach reduces computational requirements while maintaining performance during inference.
Feature Importance
@@
Figure A6: Feature importance: Using the gain attributed to each feature, the feature importance for LightGBM models are estimated for sowing (DSR) classification, left and irrigation (AWD) classification on the right.
@@
For the best performing models for DSR and AWD classification, we estimate the feature importance as shown in Fig. A6. For DSR classification, the best performing model was trained on the time period Jun 1- Sep 5 with a 4 day sampling frequency (Fig. 4, row 3). Therefore the features that start with “time_*” refer to the time period in 4 day increments beginning on Jun 1888Since we pass in both S1 timesteps and a single fixed value for Dynamic World, the Presto Encoder generates: [0] lat-lon token, [1:T+1] temporal tokens, [T+1:2T+1] Dynamic World tokens. The top two features are therefore from the time period Jun 16-20, implying that the features in those regions were most useful in classification. From Fig. A1, using the lens of sowing detection, we see that this refers to the time when the majority of PTR crops have been sowed (control/DSR has peaked and AWD has begun), which aligns with our conclusion that the sowing features are actually helpful.
@@
For AWD classification the model was trained for the period May 1 to Dec 15 with a ten day sampling frequency (Fig. 4, row 12). Similarly, the top two features (Fig. A6 left) correspond to the time periods of June 20-30, and Sep 18-28. The former is roughly the same time-period as the top-2 features for DSR, while the latter likely captures some important lag feature, since all plots are at some stage of growth in September, regardless of when they were planted.
Figure A7: Inference Workflow: Methodology followed during DSR classification predictions for 3 million polygons using S-1 timeseries. FTW (kerner2024fields) polygons are generated from bitemporal 4-bands S-2 imagery
Impact of temporal overlap:
For both DSR and AWD classification, errors occur predominantly in samples with temporally similar planting dates (Figure A5). This error pattern confirms that while our dimensional approach effectively addresses most planting date heterogeneity, classification remains challenging when practices exhibit similar temporal signatures. The elevated error rate in AWD classification in the presence of irrigation features shown in Table 2 rows 8-10 further corroborates our finding that current satellite revisit frequencies inadequately capture irrigation cycles. These rows were additionally modeled with a higher sampling frequency of 4 days, yielding even poorer results, showing that higher sampling frequency has no effect on AWD performance.
Appendix G Large-scale Inference
@@
The entire workflow for large-scale inference is outlined in Fig. A7. In order to make predictions at a plot-level, we first need to delineate agricultural plots across the state. There are no current estimates of the number of plots, although the total rice-growing area in 2021 was roughly 3.2 million hectares (ha) as surmised from Han et al. (han2022annual). Given that average plot sizes range from 0.5 to 2 ha, delineation of millions of plots would be required, making manual efforts impossible. We apply the automated method known as Fields of the World (FTW) (kerner2024fields) to delineate approximately 3 million fields across the state. The FTW model is a U-net with EfficientNet-b3 backbone trained on concatenated bitemporal 4-channel RGBN images (with and without green cover, typically a few months separated), which are then polygonized to compute object-level metrics using an IoU threshold of 0.5.
@@
Thereafter, we follow the method laid out in Fig. A3 with ascending orbit only, and use the best pretrained model for DSR classification, without any climatology inputs. The combination of pretrained models such as Presto and FTW, along with our dimensional training methodology that requires no cropping calendar information, allows us to apply this methodology to arbitrary spatial scales. For samples of the output produced over Punjab, refer to Figure A9.
Figure A8: Outlier districts for masking: The two districts shown here have very low rice plot coverage according to (han2022annual) relative to DSR mapping. The insets show both the coarse nature of the rice mask, generated at 500m resolution, as well as the potential undercounting of large areas of rice.
(a) Ludhiana district. Rural area with a canal running through the eastern end.
(b) Amristar district. Semi-rural area outside the city of Amritsar
Figure A9: FTW delineation performance samples. Polygons in white represent FTW polygons, basemap from Google Maps Satellite view. Note that FTW performs relatively well in these samples and generally for plots of this size. Overall it tends to underestimate the number of plots and the size of the plot as well as have boundary artifacts.
Validation
The task most relevant to policymaking for the PRANA program is to predict across the entire state to enable decision making at the district and sub-district level to be able to deploy resources towards increasing the adoption of DSR practices. However, the task of interpolating across roughly 3 million plots based on a model trained on just over 1,400 data points requires thorough validation. Manual validation of model estimates was beyond the scope of this project, and to produce some confidence at the district level, we aggregate the results and compare to the Government of Punjab’s DSR incentive scheme. This involves providing a lump sum of Rs. 1,500 per acre to farmers who self-report acreage under DSR. For 2024, the total reported area under DSR cultivation was 253,328 acres (progressivefarming2025). We compare these district-level estimates to han2022annual as well as our estimates in Table A1. The column “DSR Area” is the raw sum of all DSR plots extracted per district. This aggregate area is almost certainly incorrect due to pixel-level errors in plot extraction from Sentinel-2 imagery. Furthermore, we did not implement a plot-level filter for rice v/s non-rice growing plots, since our training dataset did not consist of any non-rice plots in the kharif season. Therefore, we apply the rice growing regions as a mask from Han et. al 2022, and the estimates of DSR acreage are shown under the column “DSR Area (masked)”.
The overestimation relative to government estimates reduces in the masked column, but we see a extremely large change in our DSR estimates for the districts of Sri Muktsar Sahib and Fazilka (highlighted rows in Table A1), with changes in acreage in the order of 40k. Further analysis (see Figure A8) revealed that rice growing regional estimates (han2022annual) made for 2021 are likely underestimating outcomes for 2024, given that the state of Punjab has been a rapid increase in rice cultivation areas. This the south-west region which has seen more than a doubling of rice-growing regions in the last two decades (kaur2025rice). We therefore created a hybrid version of our estimates and the masked estimates by keeping all districts masked except for Fazilka and Sri Muktsar Sahib. The comparisons of district-wise acreage across these three methods is shown in Figure A10 (left). On the right of Figure A10 we show hybrid DSR plot coverage along with the rice growing regions from (han2022annual). These two districts are also the driest regions and therefore the need to implement water-saving methods would be the highest.
For qualitative analysis of DSR plot predictions, please refer to Figure A12.
Estimator (acres $\rightarrow$)  (han2022annual) Government Estimates Our estimates
District ($\downarrow$) Rice Growing Area (2021) DSR Estimates % under DSR DSR Area % under DSR DSR Area (masked) % under DSR
Amritsar 524706 18049 3.4% 11265 2.1% 10150 1.9%
Barnala 304892 5380 1.8% 12036 3.9% 9942 3.3%
Bathinda 283112 12786 4.5% 51532 18.2% 13138 4.6%
Faridkot 229485 4330 1.9% 19852 8.7% 11283 4.9%
Fatehgarh Sahib 267082 663 0.2% 10251 3.8% 9970 3.7%
Fazilka 259095 78558 30.3% 53303 20.6% 9304 3.6%
Firozpur 410187 18452 4.5% 9287 2.3% 6554 1.6%
Gurdaspur 538681 4527 0.8% 4349 0.8% 4013 0.7%
Hoshiarpur 132811 501 0.4% 12073 9.1% 1774 1.3%
Jalandhar 398008 1673 0.4% 8683 2.2% 4764 1.2%
Kapurthala 306852 771 0.3% 4249 1.4% 2717 0.9%
Ludhiana 727297 3928 0.5% 25898 3.6% 21506 3.0%
Mansa 260315 4495 1.7% 28184 10.8% 10407 4.0%
Moga 522039 2183 0.4% 10660 2.0% 9824 1.9%
Sri Muktsar Sahib 317697 79720 25.1% 66546 20.9% 28833 9.1%
Pathankot 56211 628 1.1% 545 1.0% 105 0.2%
Patiala 679038 6267 0.9% 38061 5.6% 34677 5.1%
Rupnagar 103007 232 0.2% 3020 2.9% 2002 1.9%
Sahibzada Ajit Singh Nagar 96121 914 1.0% 3395 3.5% 1992 2.1%
Sangrur 795607 5017 0.6% 27097 3.4% 24803 3.1%
Shahid Bhagat Singh Nagar 170241 808 0.5% 7082 4.2% 4967 2.9%
Tarn Taran 531500 3446 0.6% 7236 1.4% 5758 1.1%
Table A1: Comparison of different rice growing area estimates. All areas are in acres.
Figure A10: Comparison of different masking methodologies used in this study. Left: ‘Ensemble’ refers to the DSR ensemble model outputs. ‘Ensemble Masked’ is the ensemble masked with (han2022annual). ‘Ensemble Masked Hybrid’ is where we use the outputs from Ensemble Masked except for Fazilka and Sri Muktsar Sahib. Right: The final DSR plot distributions from Ensemble Masked Hybrid overlaid on Han et al. 2022’s estimates for rice producing regions in the year 2021.
Figure A11: Examples of hand-crafted features generated from SAR backscatter timeseries for different irrigation field types. Top-left: DSR and CF. Top-right: PTR and AWD. Bottom row: PTR with CF. Left subplots show the $V$ (blue) and VH (orange) timeseries for every pixel contained within a plot, and the dark blue and dark orange show the mean (with shaded regions denoting the standard deviation) VV and VH, respectively, per plot. Right subplots show the gaussian fit to the VV/VH curve and the radar vegetation index (RVI). The dotted lines in each left-subplot denote the planting date for that plot.
(a) Ludhiana district
(b) Fazilka district
(c) Sri Muktsar Sahib district
(d) Sangrur district
Figure A12: DSR Plot predictions for different districts in Punjab, with insets showing sample crops of localized regions